OpenStudy (anonymous):

a grocery store sells an imported speciality cheesecake for $11 and its own store-baked cheesecake for$5. use the distributive property to mentally find the total cost for 6 of each type of cheesecake.

OpenStudy (anonymous):

6x11 + 6x5 is 6 x (11+5) is 6 x 16 = 96

OpenStudy (anonymous):

11*6=66 5*6=30 66+30=96